# How informative are summaries of the cosmic 21-cm signal?

### v1

@@ 3.4 Recurrent Neural Network
A Recurrent Neural Network (RNN) is a type of Neural Network (NN) architecture designed to efficiently encode sequential or time-evolving information. In our previous work (Prelogović et al., 2022), we used RNNs along the redshift axis in combination with Convolutional NNs (CNNs) along the sky-plane, to capture the anisotropic evolution of the cosmic 21-cm signal along the lightcone. Specifically, we used a long short-term memory (LSTM, Hochreiter & Schmidhuber (1997); Shi et al. (2015)) RNN, which is more stable to train compared with older RNN versions (see e.g. the review in Schmidt 2019). Using RNNs resulted in a more accurate regression of astrophysical parameters, compared to only using CNNs (e.g. Gillet et al. 2019; La Plante & Ntampaka 2019; Mangena et al. 2020; Kwon et al. 2020; Zhao et al. 2022a; Heneka 2023).
In this work, we use the flagship model from Prelogović et al. (2022), SummaryRNN, together with the exact weights from that work. The model has been trained in a supervised manner as a regressor. The database used for training had the same telescope simulator, but used an older version of the 21cmFAST simulator with a different astrophysical parametrization.
Our RNN summary in this work is the 32-dim output from the first dense layer of the network (see Table B3 in Prelogović et al. 2022, where convolutional and recurrent layers are followed by dense layers, ending with prediction of four parameters). Using such a higher-dimensional layer instead of the final parameter regressor output (e.g. Zhao et al. 2022a), makes the summary more sensitive to the general lightcone features instead of the specific model and/or parametrization used to create the training set. This is similar to so-called ?transfer learning? (e.g. Jiang et al. 2022), where one uses a NN that has been pre-trained on a different domain (i.e. database) and re-trains its last layers on a new domain. As our original RNN takes as input $25\times 25$ sky-plane slices, we roll a $25\times 25$ window over each redshift slice, averaging the corresponding dense layer outputs to obtain the final RNN summary.
3.5 Information Maximizing Neural Network
An Information Maximizing Neural Network (IMNN; Charnock et al. 2018) [^7] https://github.com/tomcharnock/IMNN is an unsupervised learning algorithm, where the NN is specifically trained to maximize the Fisher information $\det\bm{F}$ of the summary compression, at a single fixed point in parameter space $\bm{\theta}_{\text{fid}}$. IMNNs have been previously applied to galaxy LSS maps (e.g. Makinen et al. 2022) and catalog-free modeling of galaxy types (Livet et al., 2021). Here we apply them for the first time to 21-cm maps of the EoR and CD. In what follows, we outline the main ingredients of the algorithm and the training procedure.
Taking some NN architecture (in our case a simple CNN), one compresses the lightcone $\bm{l}$ into a summary vector $\bm{d}$
@@ 3.5 Information Maximizing Neural Network
The dimensionality of the summary is equal to the parameter space dimensionality ($D$). This is motivated by the fact that, for a given model parametrization, the ?score? of the real likelihood function $\nabla_{\bm{\theta}}\ln P(\bm{l}|\bm{\theta})$ is sufficient for optimal parameter recovery, and thus the dimensionality of the summary which maximizes information around one fiducial point can be the same as the dimensionality of the parameter space (e.g. Heavens et al. 2000; Charnock et al. 2018).
We train the IMMN using the following loss function (c.f. Charnock et al. 2018; Makinen et al. 2021):
@@
$L_{\mathcal{F}}({\bm{\Sigma}(\bm{\theta}_{\text{fid}})})$ is the Frobenius norm between the covariance and the unit matrix $\bm{\mathbb{1}}$, $L_{2}(W_{NN})$ is the $l_{2}$ norm of the networks’ weights. The final loss is then the sum of the negative logarithm of the Fisher information, together with the covariance and weights regularizers. The $\tanh$ term in the covariance regularizer is used to turn-off the regularization once the covariance becomes close to the unity. The main goal of the IMNN is therefore to maximize the Fisher information, while keeping the covariance of the summary fixed to unity. [^8] For any summary $\bm{d}$ of dimensionality D, one can define a new summary, $\bm{d}^{\prime}=\bm{a}+\bm{B}\bm{d}$, where $\bm{a}$ is a constant D-dim vector serving to normalize $\bm{d}^{\prime}$ to be zero mean, and $\bm{B}$ is a non-singular $D\times D$ matrix, serving to normalize $\bm{d}^{\prime}$ to be unit variance. One can then show that $\ln P(\bm{d}|\bm{\theta})=\ln P(\bm{d}^{\prime}|\bm{\theta})+\ln\det\bm{B}\,$, or equivalently $\partial\ln P(\bm{d}|\bm{\theta})/\partial\bm{\theta}_{m}=\partial\ln P(\bm{d}^{\prime}|\bm{\theta})/\partial\bm{\theta}_{m}$. This means that one always has the freedom to transform the summary $\bm{d}$ to be zero-mean, unit-variance at a single point $\bm{\theta_{\rm fid}}$, without changing the Fisher information anywhere in the parameter space. Note that the Fisher matrix is calculated only using the first term and without covariance inverse between the two derivatives (see Eq. 4). Both of these simplifications were needed in order to stabilize the training of the IMMN, and they are correct as long as the covariance is kept at unity. [^9] The original work by Charnock et al. (2018) had slightly different loss function, by having the inverse of the covariance in the Fisher term and the additional Frobenius norm between $\bm{\Sigma}^{-1}$ and $\bm{\mathbb{1}}$. We found that these choices destabilized the training and are not crucial for the convergence (see also Makinen et al. 2021).
@@
For the IMNN training, we create a database around the fiducial $\bm{\theta}_{\text{fid}}$ (see Figs. 1 and 4) for $N=M=1024$ (Eqs. 23, 24, 25), with $11264$ realizations in total. The training and validation sets are formed by splitting the database in half, following the original work of Charnock et al. (2018). We train the network using stochastic gradient descent with a batch size $b=128$, regularization strengths $\lambda=10$, $\lambda_{W}=10^{-9}$ and the Adam optimizer. To increase the effective database size and improve the training, the telescope noise is added to the simulations ?on-the-fly?. The final model has been trained on $8$ V-100 GPUs in parallel, for $\approx 120\mathrm{k}$ epochs and $\approx 20\mathrm{k}$ GPUh. The IMNN code developed for the purposes of this work we make publicly available under 21cmIMNN [^10] https://github.com/dprelogo/21cmIMNN package.
@@
Finally, we also compute the Fisher information of the concatenation of the individual 2DPS and IMNN summaries. As we shall see in the following section, these are the two best performing individual summaries. Note that one can cleanly combine summaries when performing Simulation Based Inference (SBI), as the likelihood does not need to be specified explicitly. The main restriction is the size of the required database and the stability of the SBI training (e.g. Cranmer et al. 2020; Lueckmann et al. 2021). When computing the Fisher information in this work, we are restricted to these two summaries due to the limits imposed by our database size. In particular, the minimal number of samples needed to estimate the covariance matrix is proportional to the dimensionality of the total summary (e.g. Hartlap et al. 2007). Furthermore, the dimensions of the final summary should be uncorrelated, as otherwise the covariance is singular and non-invertible. This can result in incorrect Fisher matrix estimation (see however Tegmark et al. 1997 for possible alternative solutions).
@@ 4 Database
Figure 4: Prior volume of the astrophysical parameter space. Dashed lines demarcate the fiducial parameter values to train the IMNN, 2D contours enclose 95% of the prior volume, and the points denote the 152 prior samples at which we compute the Fisher information for our summaries. Note that all units and $\log_{10}$ factors have been removed for clarity - see the main text and Fig. 1 for reference on the units.
Since we do not have a strongly-motivated fiducial choice for our galaxy parameters, we want to compare the Fisher information of 21-cm summaries across our prior volume. In order to do so, we construct a ?Fisher database?, sampling 152 parameter combinations from our prior. We use a flat prior for the X-ray emission parameters, $L_{X}$ and $E_{0}$, over the ranges shown in Fig. 4 (for more motivation of these choices see Fragos et al. 2013; Lehmer et al. 2016; Das et al. 2017). The remaining parameters characterizing star formation and ionizing emission of galaxies are constrained by observed UV luminosity functions (e.g. Bouwens et al. 2015, 2017; Oesch et al. 2018) as well as estimates of the reionization history (e.g. McGreer et al. 2015; Planck Collaboration et al. 2016). As our model is nested in the 6D Park et al. (2019) model, we use the posterior from that work, calculating the corresponding conditional distributions over $f_{*,10}$, $f_{\text{esc},10}$ and $M_{\text{turn}}$, fixing the remaining parameters to their fiducial values, as discussed in the previous section. To compute the conditional distributions, we use the conditional_kde [^11] https://github.com/dprelogo/conditional_kde code.
Our prior and samples are shown in Fig. 4. The diagonal panels illustrate the 1D prior PDFs, while the contours in the lower left enclose 95% of the 2D distributions. The 152 samples are shown as points, while the dashed vertical and horizontal lines demarcate the fiducial choice used to train our IMNN.
We simulate $N=130$ realizations for each of the 152 parameter samples, $\bm{\theta^{*}}$, in order to estimate the mean and the covariance of each summary (Eqs. 5, 6), plus an additional $1502\times 5\times M=$ 150 realizations at $\bm{\theta^{*}}\pm\delta\bm{\theta}_{m}/2$ to estimate the derivative of the mean (Eq. 10) along each of the 5 parameter basis vectors. We note that no new simulations are needed to estimate the derivative of the covariance matrix, since we use the same $L=M=15$ realizations that are used to compute the first Fisher term mentioned above (c.f. Eq. 11 and associated discussion).
@@ 5 Results
We now present the main results of this work. We first discuss the training of the IMNN in Section 5.1, then compare the performance of all of our 21-cm summaries in Section 5.2, adding a final discussion on the importance of including the second Fisher term (c.f. Eq. 4) in Section 5.3. Throughout, our figure of merit will be the PDF of the Fisher information ($\log_{10}(\det\bm{F})$), over the prior volume (see previous section). We quote both the median and the variance of this distribution. A high value of the median means that the typical parameter value is well constrained by the summary, while a narrow variance means that the constraining power of the summary is relatively constant over the prior volume. We note that the Fisher information is a measure of the average constraining power over our five parameters; we do not discuss the relative constraints on individual parameters in this work.
@@ 5.1 IMNN training
Figure 5: Fisher information of the IMNN during training. Dark (light) blue curves show the Fisher information for the training (validation) set as the IMNN is trained at $\bm{\theta}_{\rm fid}$. The violin plots show the distribution of the Fisher information over the full prior volume excluding $\bm{\theta}_{\rm fid}$, $P(\theta)$, at every interval of 10k training epochs. Black dots and error bars inside each violin denote the median and 68% C.L. of the distributions. The divergence between the Fisher information computed at $\bm{\theta}_{\rm fid}$ and its PDF over the remaining prior volume clearly indicates that the IMNN is over-specializing to the specific parameter combination, $\bm{\theta}_{\rm fid}$, used in the training.
@@
Is the compression learned by the IMNN comparably informative at other parameter combinations? We can answer this question by comparing the solid lines of Fig. 5 to the violin distributions. The latter show the distribution of the Fisher information over the full prior volume, $P(\theta)$, at every interval of 10k training epochs, excluding $\bm{\theta}_{\rm fid}$. We see that although the IMNN is not overfitting in the classical sense (the Fisher information of the validation set keeps increasing), it quickly ?over-specializes? to the specific parameter value used for training, $\bm{\theta}_{\rm fid}$. Thus the compression learned by the IMNN does not generalize well across parameter space. [^12] One solution could be to train the IMNN around several different points. However, this is extremely expensive in our application, both in terms of database size and GPU time. Furthermore, it is not clear if such procedure would result in a meaningful summary, as the concept of one point training is inherently built into the IMNN. Alternatively, it might be possible to construct a generalized IMNN, maximizing the Fisher information throughout parameter space; however, it is not obvious how this could be done. The PDF of the Fisher information over the prior volume does not change significantly after epoch 40k, with the median beginning to decrease beyond epoch 90k. We take epoch 90k as the final state of the IMNN in the remainder of this work.
@@ 5.2 Comparison of all summaries
Figure 6: Distribution of the Fisher information ($\det\bm{F}$) across the prior volume, for all summaries considered in this work. For the IMNN summary, we use the weights that give the highest median value (corresponding to 90k epochs; see Fig. 5). On the top axis we denote the corresponding relative improvement in the single parameter variance (Eq. 26), normalized to its median value of the 2DPS summary. The 2DPS has the largest median value of all of the individual summaries, while the combination of the 2DPS + IMNN dramatically outperforms every individual summary. See Table 1 for quantitative values.
@@
Table 1: Companion table to Fig. 6, summarizing the Fisher information probability distribution across the prior volume for each summary. In particular, the first column represents values of $\log_{10}(\det\bm{F})$ for the median and $68\%$ of probability volume around it. The second column shows the single parameter variance (Eq. 26), which serves as a guide to how much would the recovered variance of a parameter increase or decrease if one uses that particular summary instead of the 2DPS.
In Fig. 6 we show the main result of this work: the distribution of Fisher information across the prior volume, for all of the summary statistics we consider. On the top axis we denote the corresponding relative improvement in constraints on the single parameter variance $\sigma^{2}$. We define it as an effective variance in each dimension with no correlation, giving the same Fisher information. For a 5-dim parameter space we have:
@@
Here, one can ignore the difference in the units (i.e. scaling) between parameter dimensions, see Appendix C for more details. Specifically, the top axis of Fig. 6 denotes the single parameter variance normalized to its median value of the 2DPS summary. To give an example, a factor of $10$ times larger Fisher information for a 5-dim parameter space would result in a $37\%$ smaller single parameter variance. The median values and 68% C.L. of the Fisher information and single parameter variance are also listed in Table 1.
@@
Among the individual summaries, the 2DPS is a clear winner. In comparison, the 1DPS has roughly a factor of two smaller Fisher information across the prior volume, which translates to a $\sim 15\%$ larger single parameter variance (Eq. 26). This shows that indeed the anisotropy of the cosmic 21-cm signal helps in parameter recovery, though not very significantly. The distribution of the 2DPS Fisher information is also relatively narrow, indicating its constraining power does not vary enormously across parameter space, compared to most other summaries.
@@
The median Fisher information of the IMNN is comparable with that of the 1DPS. However, the IMNN results in the widest distribution of all of the summaries. This means that the compression that the IMNN learned at $\bm{\theta}_{\rm fid}$ can be much more informative but also much less informative, depending on the parameter choice. Therefore, if one has a good idea of where the maximum likelihood value will be (e.g. from complementary observations; Park et al. 2019; Abdurashidova et al. 2022; HERA Collaboration et al. 2023; Breitman et al. 2024), it could be beneficial to train the IMNN at that ?best guess? value and use the resulting summary in inference. An optimal strategy could be to perform inference using the 2DPS in order to obtain the maximum likelihood, train the IMNN at this ML parameter combination, then perform inference again using the resulting IMNN summary.
The wavelet summary results in a median Fisher information that is a factor of $\sim$300 smaller than that of the 2DPS at their median values (factor of $\sim$3 increase in the single parameter variance (Eq. 26). Intuitively, one would expect more information content in wavelets compared with the PS, as they effectively include higher order correlations. However, our WST followed the definition in Greig et al. (2022, 2023), in which the wavelets are only computed at one slice for each redshift chunk. Thus, the physics-rich, line-of-sight modes inside each redshift bin are lost for this choice of WST. The weaker performance of the WST compared to the power spectra implies that there is more information in the line-of-sight modes than there is in higher-order correlations of the transverse modes.
We note that recent works also using 2D wavelets achieved better results than we do here (Greig et al., 2022; Hothi et al., 2023). However, their analysis differs from ours. In particular, they use more redshift bins and they do not average the coefficients over the sky-plane (rolling the filter as described above). We confirm that without the additional smoothing that we perform in this work, the Gaussian approximation for the likelihood that is intrinsic to the Fisher estimate (c.f. Eq. 4) is notably worse, for both the wavelets and the RNN. Therefore, the higher wavelet information found by Greig et al. (2022); Hothi et al. (2023) could be partially due to a less appropriate application of the Fisher estimate. Alternatively, the information content could be improved by using 3D wavelets instead of 2D ones (see e.g. Zhao et al. 2023a).
@@
Of the individual summaries, the RNN performs the worst with a median Fisher information that is three orders of magnitude smaller than that of the 2DPS. However, its distribution is the most homogeneous across the prior volume (i.e. the narrowest). This could be interpreted as sacrificing the overall information content for robustness throughout the parameter space. Indeed the RNN was trained as a regressor over a large parameter space, and so it is understandable that it is much more consistent with respect to the IMNN which was trained on a single parameter combination. The overall low median information content of the RNN compression could be due to the fact that it was trained on a different parametrization of the cosmological signal than used in this work. We expect that retraining the RNN as a regressor on the same parametrization would improve the median information content. Indeed, Zhao et al. (2022a) find that a regressor CNN slightly outperforms the 1DPS, resulting in tighter parameter constraints when using two different parameter combinations for the mock observation. However, it is likely that this increase in the information content would come at the cost of robustness (i.e. a broader PDF), as the RNN over-specializes to features inherent in that model’s parametrization and its training set.

### v2

@@ 3.4 Recurrent neural network
A RNN is a type of NN architecture designed to efficiently encode sequential or time-evolving information. In our previous work (Prelogović et al., 2022), we used RNNs along the redshift axis in combination with convolutional neural networks (CNNs) along the sky-plane, to capture the anisotropic evolution of the cosmic 21 cm signal along the light cone. Specifically, we used a long short-term memory (Hochreiter & Schmidhuber 1997; Shi et al. 2015) RNN, which is more stable to train compared with older RNN versions (see, e.g., the review in Schmidt 2019). Using RNNs resulted in a more accurate regression of astrophysical parameters, compared to only using CNNs (e.g., Gillet et al. 2019; La Plante & Ntampaka 2019; Mangena et al. 2020; Kwon et al. 2020; Zhao et al. 2022a; Heneka 2023).
In this work, we used the flagship model from Prelogović et al. (2022), SummaryRNN, together with the exact weights from that work. The model has been trained in a supervised manner as a regressor. The database used for training had the same telescope simulator, but used an older version of the 21cmFAST simulator with a different astrophysical parametrization.
Our RNN summary in this work is the 32-dim output from the first dense layer of the network (see Table B3 in Prelogović et al. 2022, where convolutional and recurrent layers are followed by dense layers, ending with prediction of four parameters). Using such a higher-dimensional layer instead of the final parameter regressor output (e.g., Zhao et al. 2022a), makes the summary more sensitive to the general light cone features instead of the specific model and/or parametrization used to create the training set. This is similar to so-called transfer learning (e.g., Jiang et al. 2022), where one uses a NN that has been pretrained on a different domain (i.e., database) and retrains its last layers on a new domain. As our original RNN takes as input $25\times 25$ sky-plane slices, we rolled a $25\times 25$ window over each redshift slice, averaging the corresponding dense layer outputs to obtain the final RNN summary.
3.5 Information-maximizing neural network
An IMNN (Charnock et al. 2018) [^10] https://github.com/tomcharnock/IMNN is an unsupervised learning algorithm, where the NN is specifically trained to maximize the Fisher information $\det\bm{F}$ of the summary compression, at a single fixed point in parameter space $\bm{\theta}_{\text{fid}}$. IMNNs have been previously applied to galaxy large-scale structure maps (e.g., Makinen et al. 2022) and catalog-free modeling of galaxy types (Livet et al., 2021). Here we applied them for the first time to 21 cm maps of the EoR and CD. In what follows, we outline the main ingredients of the algorithm and the training procedure.
Taking some NN architecture (in our case a simple CNN), one compresses the light cone, $\bm{l}$, into a summary vector, $\bm{d}$,
@@ 3.5 Information-maximizing neural network
The dimensionality of the summary is equal to the parameter space dimensionality ($D$). This is motivated by the fact that, for a given model parametrization, the ?score? of the real likelihood function, $\nabla_{\bm{\theta}}\ln P(\bm{l}|\bm{\theta})$, is sufficient for optimal parameter recovery, and thus the dimensionality of the summary that maximizes information around one fiducial point can be the same as the dimensionality of the parameter space (e.g., Heavens et al. 2000; Charnock et al. 2018).
We trained the IMNN using the following loss function (c.f. Charnock et al. 2018; Makinen et al. 2021):
@@
$L_{\mathcal{F}}({\bm{\Sigma}(\bm{\theta}_{\text{fid}})})$ is the Frobenius norm between the covariance and the unit matrix $\bm{\mathbb{1}}$, $L_{2}(W_{NN})$ is the $l_{2}$ norm of the networks’ weights. The final loss is then the sum of the negative logarithm of the Fisher information, together with the covariance and weight regularizers. The $\tanh$ term in the covariance regularizer is used to turn-off the regularization once the covariance becomes close to the unity.
The main goal of the IMNN is therefore to maximize the Fisher information, while keeping the covariance of the summary fixed to unity. The reason we have the flexibility to normalize the covariance of the summary is the following. For any summary $\bm{d}$ of dimensionality D, one can define a new summary, $\bm{d}^{\prime}=\bm{a}+\bm{B}\bm{d}$, where $\bm{a}$ is a constant D-dim vector serving to normalize $\bm{d}^{\prime}$ to be zero mean, and $\bm{B}$ is a nonsingular $D\times D$ matrix, serving to normalize $\bm{d}^{\prime}$ to be unit variance ($\bm{B}$ is equivalent to a zero-phase component analysis (ZCA) whitening kernel; see Kessy et al. 2018). One can then show that $\ln P(\bm{d}|\bm{\theta})=\ln P(\bm{d}^{\prime}|\bm{\theta})+\ln\det\bm{B}\,$, or equivalently $\partial\ln P(\bm{d}|\bm{\theta})/\partial\bm{\theta}_{m}=\partial\ln P(\bm{d}^{\prime}|\bm{\theta})/\partial\bm{\theta}_{m}$. This means there is a freedom to transform the summary $\bm{d}$ to be zero-mean, unit-variance at a single point $\bm{\theta_{\rm fid}}$, without changing the Fisher information anywhere in the parameter space.
@@
We also note that the Fisher matrix used for computing the IMNN loss is calculated with two approximations: i) using only the first term; and ii) without the inverse covariance between the two derivatives (see Eq. 4). Both of these simplifications were needed in order to stabilize the training of the IMNN, and they are correct as long as the covariance is kept at unity. [^11] The original work by Charnock et al. (2018) had slightly different loss function, by having the inverse of the covariance in the Fisher term and the additional Frobenius norm between $\bm{\Sigma}^{-1}$ and $\bm{\mathbb{1}}$. We find that these choices destabilized the training and are not crucial for the convergence (see also Makinen et al. 2021).
@@
For the IMNN training, we created a database around the fiducial $\bm{\theta}_{\text{fid}}$ (see Figs. 1 and 5) for $N=M=1024$ (Eqs. 23, 24, 25), with $11264$ realizations in total. The training and validation sets were formed by splitting the database in half, following the original work of Charnock et al. (2018). We trained the network using stochastic gradient descent with a batch size $b=128$, regularization strengths $\lambda=10$, $\lambda_{W}=10^{-9}$ and the Adam optimizer. To increase the effective database size and improve the training, the telescope noise was added to the simulations ?on the fly.? The final model has been trained on $8$ V-100 GPUs in parallel, for $\approx 120\mathrm{k}$ epochs and $\approx 20\mathrm{k}$ GPUh. The IMNN code developed for the purposes of this work we make publicly available under 21cmIMNN [^12] https://github.com/dprelogo/21cmIMNN package.
@@
Finally, we also computed the Fisher information of the concatenation of the individual 2DPS and IMNN summaries. As we shall see in the following section, these are the two best performing individual summaries. We note that one can cleanly combine summaries when performing simulation-based inference (SBI), as the likelihood does not need to be specified explicitly. The main restriction is the size of the required database and the stability of the SBI training (e.g., Cranmer et al. 2020; Lueckmann et al. 2021). When computing the Fisher information in this work, we were restricted to these two summaries due to the limits imposed by our database size. In particular, the minimal number of samples needed to estimate the covariance matrix is proportional to the dimensionality of the total summary (e.g., Hartlap et al. 2007). Furthermore, the dimensions of the final summary should be uncorrelated, as otherwise the covariance is singular and non-invertible. This can result in incorrect Fisher matrix estimation (see however Tegmark et al. 1997 for possible alternative solutions).
@@ 4 Database
Figure 5: Prior volume of the astrophysical parameter space. Dashed lines demarcate the fiducial parameter values used to train the IMNN, 2D contours enclose 95% of the prior volume, and the points denote the 152 prior samples at which we computed the Fisher information for our summaries. Note that all units and $\log_{10}$ factors have been removed for clarity; see the main text and Fig. 1 for reference on the units.
Since we did not have a strongly motivated fiducial choice for our galaxy parameters, we wanted to compare the Fisher information of 21 cm summaries across our prior volume. To do so, we constructed a ?Fisher database,? sampling 152 parameter combinations from our prior. We used a flat prior for the X-ray emission parameters, $L_{X}$ and $E_{0}$, over the ranges shown in Fig. 5 (for more motivation of these choices see Fragos et al. 2013; Lehmer et al. 2016; Das et al. 2017). The remaining parameters characterizing star formation and ionizing emission of galaxies are constrained by observed UV luminosity functions (e.g., Bouwens et al. 2015, 2017; Oesch et al. 2018) as well as estimates of the reionization history (e.g., McGreer et al. 2015; Planck Collaboration et al. 2016). As our model is nested in the 6D Park et al. (2019) model, we used the posterior from that work, calculating the corresponding conditional distributions over $f_{*,10}$, $f_{\text{esc},10}$ and $M_{\text{turn}}$, fixing the remaining parameters to their fiducial values, as discussed in the previous section. To compute the conditional distributions, we used the conditional_kde [^13] https://github.com/dprelogo/conditional_kde code.
Our prior and samples are shown in Fig. 5. The diagonal panels illustrate the 1D prior probability distribution functions (PDFs), while the contours in the lower left enclose 95% of the 2D distributions. The 152 samples are shown as points, while the dashed vertical and horizontal lines demarcate the fiducial choice used to train our IMNN.
We simulated $N=130$ realizations for each of the 152 parameter samples, $\bm{\theta^{*}}$, in order to estimate the mean and the covariance of each summary (Eqs. 5, 6), plus an additional $1502\times 5\times M=$ 150 realizations at $\bm{\theta^{*}}\pm\delta\bm{\theta}_{m}/2$ to estimate the derivative of the mean (Eq. 10) along each of the 5 parameter basis vectors. We note that no new simulations are needed to estimate the derivative of the covariance matrix, since we used the same $L=M=15$ realizations that were used to compute the first Fisher term mentioned above (c.f. Eq. 11 and associated discussion). In Appendix D we show convergence tests of these choices for estimating the Fisher information of our summaries. We caution that the WST and RNN summaries require a larger dataset to fully converge.
@@ 5 Results
We now present the main results of this work. We first discuss the training of the IMNN in Section 5.1, then compare the performance of all of our 21 cm summaries in Section 5.2, adding a final discussion on the importance of including the second Fisher term (c.f. Eq. 4) in Section 5.3. Throughout, our figure of merit will be the PDF of the Fisher information ($\log_{10}(\det\bm{F})$), over the prior volume (see Sect. 4). We quote both the median and the variance of this distribution. A high value of the median means that the typical parameter value is well constrained by the summary, while a narrow variance means that the constraining power of the summary is relatively constant over the prior volume. We note that the Fisher information is a measure of the average constraining power over our five parameters; we do not discuss the relative constraints on individual parameters in this work.
@@ 5.1 IMNN training
Figure 6: Fisher information of the IMNN during training. Dark (light) blue curves show the Fisher information for the training (validation) set of the IMNN trained at $\bm{\theta}_{\rm fid}$. The violin plots show the distribution of the Fisher information over the full prior volume excluding $\bm{\theta}_{\rm fid}$, $P(\theta)$, at every interval of the 10k training epochs. Black dots and error bars inside each violin denote the median and 68% CLs of the distributions. The divergence between the Fisher information computed at $\bm{\theta}_{\rm fid}$ and its PDF over the remaining prior volume clearly indicates that the IMNN is over-specializing to the specific parameter combination, $\bm{\theta}_{\rm fid}$, used in the training.
@@
Is the compression learned by the IMNN comparably informative at other parameter combinations? We can answer this question by comparing the solid lines of Fig. 6 to the violin distributions. The latter show the distribution of the Fisher information over the full prior volume, $P(\theta)$, at every interval of 10k training epochs, excluding $\bm{\theta}_{\rm fid}$. We see that although the IMNN is not overfitting in the classical sense (the Fisher information of the validation set keeps increasing), it quickly ?over-specializes? to the specific parameter value used for training, $\bm{\theta}_{\rm fid}$. Thus, the compression learned by the IMNN does not generalize well across parameter space. [^14] One solution could be to train the IMNN around several different points. However, this is extremely expensive in our application, both in terms of database size and GPU time. Furthermore, it is not clear if such a procedure would result in a meaningful summary, as the concept of one point training is inherently built into the IMNN. Alternatively, it might be possible to construct a generalized IMNN, maximizing the Fisher information throughout parameter space; however, it is not obvious how this could be done. The PDF of the Fisher information over the prior volume does not change significantly after epoch 40k, with the median beginning to decrease beyond epoch 90k. We took epoch 90k as the final state of the IMNN in the remainder of this work.
@@ 5.2 Comparison of all summaries
Figure 7: Distribution of the Fisher information ($\det\bm{F}$) across the prior volume, for all summaries considered in this work. For the IMNN summary, we used the weights that give the highest median value (corresponding to 90k epochs; see Fig. 6). On the top axis we denote the corresponding relative improvement in the single parameter variance (Eq. 26), normalized to its median value of the 2DPS summary. The 2DPS has the largest median value of all of the individual summaries, while the combination of the 2DPS + IMNN dramatically outperforms every individual summary. See Table 1 for quantitative values.
Table 1: Distributions of the Fisher information.
@@
 [^15] Companion table to Fig. 7, summarizing the Fisher information probability distribution across the prior volume for each summary. In particular, the first column represents values of $\log_{10}(\det\bm{F})$ for the median and $68\%$ of probability volume around it. The second column shows the single parameter variance (Eq. 26), which serves as a guide to how much would the recovered variance of a parameter increase or decrease if one uses that particular summary instead of the 2DPS.
In Fig. 7 we show the main result of this work: the distribution of Fisher information across the prior volume, for all of the summary statistics we consider. On the top axis we denote the corresponding relative improvement in constraints on the single parameter variance $\sigma^{2}$. We define it as an effective variance in each dimension with no correlation, giving the same Fisher information. For a 5D parameter space we have
@@
Here, one can ignore the difference in the units (i.e., scaling) between parameter dimensions (see Appendix C for more details). Specifically, the top axis of Fig. 7 denotes the single parameter variance normalized to its median value of the 2DPS summary. To give an example, a factor of $10$ times larger Fisher information for a 5-dim parameter space would result in a $37\%$ smaller single parameter variance. The median values and 68% confidence levels (CLs) of the Fisher information and single parameter variance are also listed in Table 1.
@@
Among the individual summaries, the 2DPS is a clear winner. In comparison, the 1DPS has roughly a factor of 2 smaller Fisher information across the prior volume, which translates to a $\sim 15\%$ larger single parameter variance (Eq. 26). This difference between the 1DPS and 2DPS shows that indeed the anisotropy of the cosmic 21 cm signal helps in parameter recovery, though not very significantly. The distribution of the 2DPS Fisher information is also relatively narrow, indicating its constraining power does not vary enormously across parameter space, compared to most other summaries.
@@
The median Fisher information of the IMNN is comparable with that of the 1DPS. However, the IMNN results in the widest distribution of all of the summaries. This means that the compression that the IMNN learned at $\bm{\theta}_{\rm fid}$ can be much more informative but also much less informative, depending on the parameter choice. Therefore, if one has a good idea of where the maximum likelihood value will be (e.g., from complementary observations; Park et al. 2019; Abdurashidova et al. 2022; HERA Collaboration et al. 2023; Breitman et al. 2024), it could be beneficial to train the IMNN at that best guess value and use the resulting summary in inference. An optimal strategy could be to perform inference using the 2DPS in order to obtain the maximum likelihood estimate (MLE), train the IMNN at this MLE parameter combination, then perform inference again using the resulting IMNN summary.
@@
The RNN performs worse than the two PS summaries, resulting in a factor of 2 larger single parameter variance compared to the 2DPS. However, its distribution is the most homogeneous across the prior volume (i.e., the narrowest). This could be interpreted as sacrificing the overall information content for robustness throughout the parameter space. The RNN was trained as a regressor over a large parameter space, and so it is understandable that it is much more consistent with respect to the IMNN, which was trained on a single parameter combination. The overall low median information content of the RNN compression could be due to the fact that it was trained on a different parametrization of the cosmological signal than used in this work. We expect that retraining the RNN as a regressor on the same parametrization would improve the median information content. Indeed, Zhao et al. (2022a) find that a regressor CNN slightly outperforms the 1DPS, resulting in tighter parameter constraints when using two different parameter combinations for the mock observation. However, it is likely that this increase in the information content would come at the cost of robustness (i.e., a broader PDF), as the RNN over-specializes to features inherent in that model’s parametrization and its training set. Additionally, we caution the Fisher estimate for the RNN summary needs a larger dataset to fully converge (c.f. Appendix D).
The wavelet summary results in a median Fisher information that is a factor of $\sim$300 smaller than that of the 2DPS at their median values (factor of $\sim$3 increase in the single parameter variance (Eq. 26). Intuitively, one would expect more information content in wavelets compared with the PS, as they effectively include higher-order correlations. However, our WST followed the definition in Greig et al. (2022, 2023), in which the wavelets are only computed at one slice for each redshift chunk. Thus, the physics-rich, line-of-sight modes inside each redshift bin are lost for this choice of WST. The weaker performance of the WST (as defined in this work) compared to the power spectra implies that there is more information in the line-of-sight modes than there is in higher-order correlations of the transverse modes.
We note that recent works also using 2D wavelets achieved better results than we do here (Greig et al., 2022; Hothi et al., 2023). However, their analysis differs from ours. In particular, they use more redshift bins and they do not average the coefficients over the sky-plane (rolling the filter as described above). We confirm that without the additional smoothing that we performed in this work, the Gaussian approximation for the likelihood that is intrinsic to the Fisher estimate (c.f. Eq. 4) is notably worse, for both the wavelets and the RNN. Therefore, the higher wavelet information found by Greig et al. (2022); Hothi et al. (2023) could be partially due to a less appropriate application of the Fisher estimate. Alternatively, the information content could be improved by using 3D wavelets instead of 2D ones (see, e.g., Zhao et al. 2023a). As for the RNN, we caution that the WST needs a larger dataset for the Fisher to fully converge (c.f. Appendix D).